# SurfSLAM: Sim-to-Real Underwater Stereo Reconstruction For Real-Time SLAM

### v1

@@ Abstract
Localization and mapping are core perceptual capabilities for underwater robots. Stereo cameras provide a low-cost means of directly estimating metric depth to support these tasks. However, despite recent advances in stereo depth estimation on land, computing depth from image pairs in underwater scenes remains challenging. In underwater environments, images are degraded by light attenuation, visual artifacts, and dynamic lighting conditions. Furthermore, real-world underwater scenes frequently lack rich texture useful for stereo depth estimation and 3D reconstruction. As a result, stereo estimation networks trained on in-air data cannot transfer directly to the underwater domain. In addition, there is a lack of real-world underwater stereo datasets for supervised training of neural networks. Poor underwater depth estimation is compounded in stereo-based Simultaneous Localization and Mapping (SLAM) algorithms, making it a fundamental challenge for underwater robot perception. To address these challenges, we propose a novel framework that enables sim-to-real training of underwater stereo disparity estimation networks using simulated data and self-supervised finetuning. We leverage our learned depth predictions to develop SurfSLAM, a novel framework for real-time underwater SLAM that fuses stereo cameras with IMU, barometric, and Doppler Velocity Log (DVL) measurements. Lastly, we collect a challenging real-world dataset of shipwreck surveys using an underwater robot. Our dataset features over 24,000 stereo pairs, along with high-quality, dense photogrammetry models and reference trajectories for evaluation. Through extensive experiments, we demonstrate the advantages of the proposed training approach on real-world data for improving stereo estimation in the underwater domain and for enabling accurate trajectory estimation and 3D reconstruction of complex shipwreck sites. Upon publication, code, real-world underwater data, and simulated training data will be made publicly available at https://umfieldrobotics.github.io/SurfSLAM/.
@@ II.A Visual SLAM
II.B Underwater Visual SLAM
@@
II.C Stereo Depth Estimation
@@
II.D Deep Learning for Underwater Stereo
@@ III Method Overview
The discussion of the proposed technical approach is organized as follows: Subsection III.A provides the notation and coordinate conventions used throughout the paper. Section IV presents our proposed methodology for sim-to-real training of underwater stereo disparity estimation. Then, Section V describes our SurfSLAM system, which fuses stereo depth estimation with acoustic and inertial sensing for real-time localization and dense mapping of complex underwater scenes.
III.A Preliminaries
III.A.1 Notation
@@ III.A.2 Frame Definitions
IV Underwater Stereo Estimation
@@
IV.A Simulated Data Generation
@@
IV.B Underwater Data Augmentation
@@
IV.B.1 Water Column Effects
@@ IV.B.5 Particles
IV.B.6 Full Augmentation Pipeline
@@ IV.C.1 Loss Functions
IV.C.2 Direct Disparity Supervision
@@
IV.C.3 Self-Supervised Warping Loss
@@ IV.C.4 Occam Regularizer
IV.C.5 Total Loss Function
@@ V.B Measurement Definitions
V.C Acoustic-Inertial Pose Graph Formulation
@@ V.D Global Registration
V.D.1 Global Place Recognition
@@
V.D.2 Visual Feature Matching
@@
V.D.3 Geometric Validation
Using the stereo estimation described in IV, a depth image is computed for both the query and matched images. Using these depth images, the filtered 2D correspondences estimated by SuperGlue are lifted to 3D. These 3D feature correspondences are used to estimate a 3D transformation, $\mathbf{T}_{qm}^{C}$, between the two keyframes. This is again performed using RANSAC to reject poor correspondences. If insufficient inliers are detected, the registration is rejected. Otherwise, the RANSAC-estimated transform is refined using GICP [59] on the full point clouds estimated by the stereo depth estimation network, yielding a final geometric relative pose measurement $\tilde{\mathbf{T}}_{qm}^{C}\in\mathrm{SE}(3)$, which we simply denote as the stereo measurement, $\mathcal{S}_{qm}$.
@@ V.D.3 Geometric Validation
V.D.4 Pose Graph Formulation
We incorporate $\mathcal{S}_{qm}$ as a factor added to the acoustic-inertial tracking to constrain the relative motion between the keyframes. Before adding the registration to the pose graph, a final probabilistic test is performed. A stereo factor, $\mathcal{S}_{qm}$, is added to the graph after keyframes at time $t_{q}$ and $t_{m}$ have been optimized using the acoustic-inertial factors described in Section V.B. As a result, the tracking system can be queried for a belief over $T^{C}_{qm}$, including mean $\mu_{T^{C}_{qm}}$ and marginal covariance $\Sigma_{T^{C}_{qm}}$. We compute the Mahalanobis distance between the stereo registration result and the tracker’s estimate according to
@@ V.E Mapping
VI Experiments
VI.A Data Collection and Generation
@@
Our collected dataset, which we call Stereo Underwater Dataset for Shipwrecks (SUDS) contains three subsets: (i) stereo evaluation, (ii) stereo training, and (iii) SLAM evaluation. Our simulation dataset, which we call UnderWater Simulation (UWSim) contains simulated data of realistic underwater scenes generated as described in Section IV.A, and is used for training our stereo networks. This differs from alternative datasets used in underwater stereo training, which use either simulated images from non-underwater images or lack the visual properties of underwater sequences.
@@
VI.A.1 Real-world Stereo Dataset
The real-world stereo training dataset consists of twenty-two real-world sequences without accompanying ground-truth data, intended for self-supervision, with detailed statistics in Table 1. This dataset is composed of a variety of underwater lighting conditions, turbidity levels, and distractors. In addition to the stereo images, this dataset also consists of the stereo intrinsic and extrinsic parameters, making the dataset suitable for self-supervised training, as described in Section IV.C.3.
@@ VI.A.1 Real-world Stereo Dataset
VI.A.2 Real-world SLAM Datasets
@@
Ground truth is generated by taking a downsampled 5 fps version of the 1080p stereo video and running the metric photogrammetry pipeline, detailed in Section VI.B.1. This ensured that even in regions where there is low texture, the photogrammetry would track the camera poses and yield a near-full trajectory of the robot with minimal dropouts. To the extent of our knowledge, the SLAM sequence of the SUDS dataset presents the first publicly released in-the-wild dataset with ground truth trajectory and maps for a robot equipped with an IMU, DVL, barometer, and stereo camera. We refer to these three scenes as Boiler, Engine, and Long. A reconstruction of the Boiler and Long scenes, along with the ground truth camera trajectories are shown in the second row of Figure 6.
@@
VI.B Generating Ground Truth for Real Data
@@
VI.B.1 Metric Photogrammetry Pipeline
@@
VI.B.2 Stereo-IMU SLAM Metric Optimization
@@
VI.B.3 Stereo-Only Stereo Metric Optimization
@@
VI.C Stereo Training Details
@@ VI.D.1 Stereo Baselines
Table 3: We separate stereo evaluations into Large Models with ViT-L backbones and Small Models that use ViT-S or no vision transformer. In both cases, ‘Ours’ refers to DEFOM with our simulation and self-supervised training. EPE (End Point Error) is measured in pixels. BP-1.0 and D1 are described in Section VI.E.1. Lower is better for all metrics.
@@ VI.D.2 SLAM Baselines
VI.E Evaluation Metrics
VI.E.1 Disparity Metrics
@@ VI.F Computation Details
VII Stereo Evaluation Results
@@
VII.A Ablation Study on Stereo Depth Estimation
@@
VIII SLAM Evaluation Results
@@
VIII.A Trajectory Tracking Evaluation
@@ VIII.B Mapping Evaluation
VIII.C Compute Comparison and Runtime

### v2

@@ Abstract
Localization and mapping are core perceptual capabilities for underwater robots. Stereo cameras provide a low-cost means of directly estimating metric depth to support these tasks. However, despite recent advances in stereo depth estimation on land, computing depth from image pairs in underwater scenes remains challenging. In underwater environments, images are degraded by light attenuation, visual artifacts, and dynamic lighting conditions. Furthermore, real-world underwater scenes frequently lack rich texture useful for stereo depth estimation and 3D reconstruction. As a result, stereo estimation networks trained on in-air data cannot transfer directly to the underwater domain. In addition, there is a lack of real-world underwater stereo datasets for supervised training of neural networks. Poor underwater depth estimation is compounded in stereo-based Simultaneous Localization and Mapping (SLAM) algorithms, making it a fundamental challenge for underwater robot perception. To address these challenges, we propose a novel framework that enables sim-to-real training of underwater stereo disparity estimation networks using simulated data and self-supervised finetuning. We leverage our learned depth predictions to develop SurfSLAM, a novel framework for real-time underwater SLAM that fuses stereo cameras with IMU, barometric, and Doppler Velocity Log (DVL) measurements. Lastly, we collect a challenging real-world dataset of shipwreck surveys using an underwater robot. Our dataset features over 24,000 stereo pairs, along with high-quality, dense photogrammetry models and reference trajectories for evaluation. Through extensive experiments, we demonstrate the advantages of the proposed training approach on real-world data for improving stereo estimation in the underwater domain and for enabling accurate trajectory estimation and 3D reconstruction of complex shipwreck sites. Code, real-world underwater data, and simulated training data will be made publicly available.
@@ 2.1 Visual SLAM
2.2 Underwater Visual SLAM
@@
2.3 Stereo Depth Estimation
@@
2.4 Deep Learning for Underwater Stereo
@@ 3 Method Overview
The discussion of the proposed technical approach is organized as follows: Subsection 3.1 provides the notation and coordinate conventions used throughout the paper. Section 4 presents our proposed methodology for sim-to-real training of underwater stereo disparity estimation. Then, Section 5 describes our SurfSLAM system, which fuses stereo depth estimation with acoustic and inertial sensing for real-time localization and dense mapping of complex underwater scenes.
3.1 Preliminaries
3.1.1 Notation
@@ 3.1.2 Frame Definitions
4 Underwater Stereo Estimation
@@
4.1 Simulated Data Generation
@@
4.2 Underwater Data Augmentation
@@
4.2.1 Water Column Effects
@@ 4.2.5 Particles
4.2.6 Full Augmentation Pipeline
@@ 4.3.1 Loss Functions
4.3.2 Direct Disparity Supervision
@@
4.3.3 Self-Supervised Warping Loss
@@ 4.3.4 Occam Regularizer
4.3.5 Total Loss Function
@@ 5.2 Measurement Definitions
5.3 Acoustic-Inertial Pose Graph Formulation
@@ 5.4 Global Registration
5.4.1 Global Place Recognition
@@
5.4.2 Visual Feature Matching
@@
5.4.3 Geometric Validation
Using the stereo estimation described in 4, a depth image is computed for both the query and matched images. Using these depth images, the filtered 2D correspondences estimated by SuperGlue are lifted to 3D. These 3D feature correspondences are used to estimate a 3D transformation, $\mathbf{T}_{qm}^{C}$, between the two keyframes. This is again performed using RANSAC to reject poor correspondences. If insufficient inliers are detected, the registration is rejected. Otherwise, the RANSAC-estimated transform is refined using GICP [59] on the full point clouds estimated by the stereo depth estimation network, yielding a final geometric relative pose measurement $\tilde{\mathbf{T}}_{qm}^{C}\in\mathrm{SE}(3)$, which we simply denote as the stereo measurement, $\mathcal{S}_{qm}$.
@@ 5.4.3 Geometric Validation
5.4.4 Pose Graph Formulation
We incorporate $\mathcal{S}_{qm}$ as a factor added to the acoustic-inertial tracking to constrain the relative motion between the keyframes. Before adding the registration to the pose graph, a final probabilistic test is performed. A stereo factor, $\mathcal{S}_{qm}$, is added to the graph after keyframes at time $t_{q}$ and $t_{m}$ have been optimized using the acoustic-inertial factors described in Section 5.2. As a result, the tracking system can be queried for a belief over $T^{C}_{qm}$, including mean $\mu_{T^{C}_{qm}}$ and marginal covariance $\Sigma_{T^{C}_{qm}}$. We compute the Mahalanobis distance between the stereo registration result and the tracker’s estimate according to
@@ 5.5 Mapping
6 Experiments
6.1 Data Collection and Generation
@@
Our collected dataset, which we call Stereo Underwater Dataset for Shipwrecks (SUDS) contains three subsets: (i) stereo evaluation, (ii) stereo training, and (iii) SLAM evaluation. Our simulation dataset, which we call UnderWater Simulation (UWSim) contains simulated data of realistic underwater scenes generated as described in Section 4.1, and is used for training our stereo networks. This differs from alternative datasets used in underwater stereo training, which use either simulated images from non-underwater images or lack the visual properties of underwater sequences.
@@
6.1.1 Real-world Stereo Dataset
The real-world stereo training dataset consists of twenty-two real-world sequences without accompanying ground-truth data, intended for self-supervision, with detailed statistics in Table 1. This dataset is composed of a variety of underwater lighting conditions, turbidity levels, and distractors. In addition to the stereo images, this dataset also consists of the stereo intrinsic and extrinsic parameters, making the dataset suitable for self-supervised training, as described in Section 4.3.3.
@@ 6.1.1 Real-world Stereo Dataset
6.1.2 Real-world SLAM Datasets
@@
Ground truth is generated by taking a downsampled 5 fps version of the 1080p stereo video and running the metric photogrammetry pipeline, detailed in Section 6.2.1. This ensured that even in regions where there is low texture, the photogrammetry would track the camera poses and yield a near-full trajectory of the robot with minimal dropouts. To the extent of our knowledge, the SLAM sequence of the SUDS dataset presents the first publicly released in-the-wild dataset with ground truth trajectory and maps for a robot equipped with an IMU, DVL, barometer, and stereo camera. We refer to these three scenes as Boiler, Engine, and Long. A reconstruction of the Boiler and Long scenes, along with the ground truth camera trajectories are shown in the second row of Figure 6.
@@
6.2 Generating Ground Truth for Real Data
@@
6.2.1 Metric Photogrammetry Pipeline
@@
6.2.2 Stereo-IMU SLAM Metric Optimization
@@
6.2.3 Stereo-Only Stereo Metric Optimization
@@
6.3 Stereo Training Details
@@ 6.4.1 Stereo Baselines
Table 3: We separate stereo evaluations into Large Models with ViT-L backbones and Small Models that use ViT-S or no vision transformer. In both cases, ‘Ours’ refers to DEFOM with our simulation and self-supervised training. EPE (End Point Error) is measured in pixels. BP-1.0 and D1 are described in Section 6.5.1. Lower is better for all metrics.
@@ 6.4.2 SLAM Baselines
6.5 Evaluation Metrics
6.5.1 Disparity Metrics
@@ 6.6 Computation Details
7 Stereo Evaluation Results
@@
7.1 Ablation Study on Stereo Depth Estimation
@@
8 SLAM Evaluation Results
@@
8.1 Trajectory Tracking Evaluation
@@ 8.2 Mapping Evaluation
8.3 Compute Comparison and Runtime